# Vector valued estimates for matrix weighted maximal operators and product $\mathrm{BMO}$

### v2

@@ 2.6 Matrix $A_{p}$ weights
2.6.1 One-parameter matrix $A_{p}$ weights
@@
Let $W$ be a $(d\times d)$ matrix valued weight on $\mathbb{R}^{n}$. We say that $W$ is a one-parameter $d\times d$-matrix valued $A_{p}$ weight if
$[W]_{A_{p}(\mathbb{R}^{n})}\coloneq\sup_{Q}\frac{1}{|Q|}\int_{Q}\left(\frac{1}{|Q|}\int_{Q}|W(x)^{1/p}W(y)^{-1/p}|^{p^{\prime}}\,\mathrm{dm}(y)\right)^{p/p^{\prime}}\,\mathrm{dm}(x)<\infty,$
where the supremum is taken over all cubes $Q$ in $\mathbb{R}^{n}$. Note that if $W$ is a $(d\times d)$ matrix valued $A_{p}$ weight on $\mathbb{R}^{n}$, then $W^{\prime}\coloneq W^{-1/(p-1)}$ is a $(d\times d)$ matrix valued $A_{p^{\prime}}$ weight on $\mathbb{R}^{n}$ with $[W^{\prime}]_{A_{p^{\prime}}(\mathbb{R}^{n})}^{1/p^{\prime}}\sim_{p,d}[W]_{A_{p}(\mathbb{R}^{n})}^{1/p}$, and
$[W]_{A_{p}(\mathbb{R}^{n})}\sim_{p,d}\sup_{Q}|\mathcal{W}^{\prime}_{Q}\mathcal{W}_{Q}|^{p},$
where the reducing matrices for $W$ correspond to exponent $p$, and those for $W^{\prime}$ correspond to exponent $p^{\prime}$.
If $\mathcal{D}$ is any dyadic grid in $\mathbb{R}^{n}$, we define
$[W]_{A_{p},\mathcal{D}}\coloneq\sup_{Q\in\mathcal{D}}\frac{1}{|Q|}\int_{Q}\left(\frac{1}{|Q|}\int_{Q}|W(x)^{1/p}W(y)^{-1/p}|^{p^{\prime}}\,\mathrm{dm}(y)\right)^{p/p^{\prime}}\,\mathrm{dm}(x),$
and we say that $W$ is a one-parameter $(d\times d)$ matrix valued $\mathcal{D}$-dyadic $A_{p}$ weight if $[W]_{A_{p},\mathcal{D}}<\infty$.
2.6.2 Biparameter matrix $A_{p}$ weights
Let $W$ be a $(d\times d)$ matrix valued weight on $\mathbb{R}^{n}\times\mathbb{R}^{m}$. We say that $W$ is a biparameter $(d\times d)$ matrix valued $A_{p}$ weight if
$[W]_{A_{p}(\mathbb{R}^{n}\times\mathbb{R}^{m})}\coloneq\sup_{R}\frac{1}{|R|}\int_{R}\left(\frac{1}{|R|}\int_{R}|W(x)^{1/p}W(y)^{-1/p}|^{p^{\prime}}\,\mathrm{dm}(y)\right)^{p/p^{\prime}}\,\mathrm{dm}(x)<\infty,$
@@ 2.6.2 Biparameter matrix $A_{p}$ weights
where the supremum is taken over all rectangles $R$ in $\mathbb{R}^{n}$ (with sides parallel to the coordinate axes). Note that if $W$ is a $(d\times d)$ matrix valued biparameter $A_{p}$ weight on $\mathbb{R}^{n}\times\mathbb{R}^{m}$, then $W^{\prime}\coloneq W^{-1/(p-1)}$ is a $(d\times d)$ matrix valued biparameter $A_{p^{\prime}}$ weight on $\mathbb{R}^{n}\times\mathbb{R}^{m}$ with $[W^{\prime}]_{A_{p^{\prime}}(\mathbb{R}^{n}\times\mathbb{R}^{m})}^{1/p^{\prime}}\sim_{d,p}[W]_{A_{p}(\mathbb{R}^{n}\times\mathbb{R}^{m})}^{1/p}$, and
$[W]_{A_{p}(\mathbb{R}^{n}\times\mathbb{R}^{m})}\sim_{p,d}\sup_{R}|\mathcal{W}^{\prime}_{R}\mathcal{W}_{R}|^{p},$
where the reducing matrices for $W$ correspond to exponent $p$, and those for $W^{\prime}$ correspond to exponent $p^{\prime}$.
If $\bm{\mathcal{D}}$ is any product dyadic grid in $\mathbb{R}^{n}\times\mathbb{R}^{m}$, we define
$[W]_{A_{p},\bm{\mathcal{D}}}\coloneq\sup_{R\in\bm{\mathcal{D}}}\frac{1}{|R|}\int_{R}\left(\frac{1}{|R|}\int_{R}|W(x)^{1/p}W(y)^{-1/p}|^{p^{\prime}}\,\mathrm{dm}(y)\right)^{p/p^{\prime}}\,\mathrm{dm}(x);$
we say that $W$ is a biparameter $(d\times d)$ matrix valued $\bm{\mathcal{D}}$-dyadic $A_{p}$ weight if $[W]_{A_{p},\bm{\mathcal{D}}}<\infty$.
2.6.3 One-parameter restriction of biparameter matrix $A_{p}$ weights
Let $1<p<\infty.$ Let $W$ be a $(d\times d)$ matrix biparameter $A_{p}$ weight on $\mathbb{R}^{n}\times\mathbb{R}^{m}.$ For a.e. $x_{1}\in\mathbb{R}^{n},$ set $W_{x_{1}}(x_{2})\coloneq W(x_{1},x_{2}),$ $x_{2}\in\mathbb{R}^{m}.$ It is proved in [DKPS2024] that
$[W_{x_{1}}]_{A_{p}(\mathbb{R}^{m})}\lesssim_{p,d}[W]_{A_{p}(\mathbb{R}^{n}\times\mathbb{R}^{m})}.$
Of course, the dyadic version of this is also true. Moreover, both versions remain true if one “slices” with respect to the first coordinate instead of the second one.
Fix any cube $Q$ in $\mathbb{R}^{m}.$ For a.e. $x_{1}\in\mathbb{R}^{n},$ let $\mathcal{W}_{x_{1},Q}$ be the reducing operator of $W_{x_{1}}(x_{2})\coloneq W(x_{1},x_{2}),$ $x_{2}\in\mathbb{R}^{m}$ over $Q$ with respect to the exponent $p.$ Set $W_{Q}(x_{1})\coloneq\mathcal{W}_{x_{1},Q}^{p},$ for a.e. $x_{1}\in\mathbb{R}^{n}.$ It is shown in [DKPS2024] that
$[W_{Q}]_{A_{p}(\mathbb{R}^{n})}\lesssim_{p,d}[W]_{A_{p}(\mathbb{R}^{n}\times\mathbb{R}^{m})}.$ (2.4)
Of course, the dyadic version of this is also true. Moreover, both versions remain true if one “slices” with respect to the first coordinate instead of the second one.
@@ 2.7 Convex body valued functions
Here we consider functions taking values in the collection of closed bounded symmetric convex sets of $\mathbb{C}^{d}.$ Take into account that by symmetric here we mean complex symmetric. That is, a set $A\subseteq\mathbb{C}^{d}$ is complex symmetric (or just symmetric in this article) if for every $u\in A$ and every $\lambda\in\mathbb{C}$ with $|\lambda|=1$ it is also the case that $\lambda u\in A.$ Of course, if the set $A$ is convex in addition to symmetric, it will also be the case that, for every $u\in A$ and every $\lambda\in\mathbb{C}$ with $|\lambda|\leq 1,$ also $\lambda u\in A.$ In other words, the “symmetric convex sets” as we have defined them are precisely the balanced convex sets. We will denote the set of closed subsets of $\mathbb{C}^{d}$ by $\mathcal{K}(\mathbb{C}^{d}),$ or just $\mathcal{K}$ when the dimension of the ambient space is clear by the context. In addition, we define a convex body to be a closed bounded convex and symmetric subset of $\mathbb{C}^{d}.$ We will use the symbol $\mathcal{K}_{\mathrm{bcs}}(\mathbb{C}^{d})$ to denote the set of convex bodies on $\mathbb{C}^{d}$ and, whenever the dimension of the ambient space is unambiguous, simply by $\mathcal{K}_{\mathrm{bcs}}.$ We focus now on functions $F\colon\mathbb{R}^{n}\rightarrow\mathcal{K}_{\mathrm{bcs}}(\mathbb{C}^{d})$ and we gather the definitions and basic properties that we will need for our results. Some of these definitions can be found in more general forms in the texts that we cite, we will restrict though to the cases that are of interest to us to avoid an excess of concepts. For an introduction to such functions, their properties and how to define their integrals see [Cruz_Uribe_Bownik_Extrapolation, Cruz_Uribe_Extrapolation] and for a detailed exposition see [Aubin_2009].
Since through the article we will consider both functions taking values in $\mathbb{C}^{d}$ and functions taking values in $\mathcal{K}_{\mathrm{bcs}},$ we will use a typographic convention to avoid confusion. We will denote functions taking values in $\mathbb{C}^{d}$ with lowercase letters $f,g,h,\ldots$ On the other hand, we will use uppercase letters $F,G,H,\ldots$ to denote functions taking values in $\mathcal{K}_{\mathrm{bcs}}(\mathbb{C}^{d}).$ In any case, we will also explicitly state the target space of the functions we use.
@@
for every $x\in\mathbb{R}^{n}$ (see [Aubin_2009, Theorem 8.1.4]). The functions that form such sequences will be called selection functions for $F.$ In addition, the set of all selection functions for $F$ is denoted by $S^{0}(F).$ Observe that this is the same as the set of all measurable functions $f$ such that $f(x)\in F(x)$ for every $x\in\mathbb{R}^{n}.$ Here we restrict to functions taking values on $\mathcal{K}_{\mathrm{bcs}}(\mathbb{C}^{d})$ due to the applications that we consider later. Nonetheless, the previous definitions and concepts apply verbatim to functions taking values on $\mathcal{K}(\mathbb{C}^{d}).$ When one is interested in the norm of $F,$ it is possible to restrict to selection functions, as the following lemma shows (see [Cruz_Uribe_Bownik_Extrapolation, Lemma 3.9]). We include its proof for completeness, although the arguments of Bownik and Cruz-Uribe are equally valid for complex convex bodies in this case.
@@
$F_{0}(x)=\{v\in F(x)\colon|v|=g_{0}(x)\}=F(x)\cap\mathbf{{S}}_{\mathbb{C}^{{d}}}$
taking values on $\mathcal{K}(\mathbb{C}^{d}),$ where $\mathbf{{S}}_{\mathbb{C}^{{d}}}=\{v\in\mathbb{C}^{d}\colon|v|=1\}$ denotes the complex $(d-1)$-dimensional sphere. Since both $F$ and $g_{0}\mathbf{{S}}_{\mathbb{C}^{{d}}}$ are measurable, so is $F_{0}$ (see [Aubin_2009, Theorems 8.2.2 and 8.2.4]).
@@
Given a $(d\times d)$-matrix weight $W,$ we define the weighted Lebesgue space $\mathrm{L}^{p}(W)$ of convex body valued functions as the set of functions $F$ for which $W(x)^{1/p}F(x)\in\mathrm{L}^{p}.$
@@ 3 Vector valued extensions of operators on convex body valued functions
The aim of this section is to prove vector valued estimates for convex body valued functions with matrix weights, as well as to apply it to vector valued estimates for the matrix weighted maximal operators for vector valued functions (see Subsection 3.1 for the definitions and details). This will be a consequence of an extrapolation result on spaces of functions with values on $\mathcal{K}_{\mathrm{bcs}}$ and it is a minor modification of the analogous result for matrix weights and vector valued functions due to Bownik and Cruz-Uribe (see [Cruz_Uribe_Bownik_Extrapolation, Section 9]). Before stating the extrapolation theorem, we introduce the concept of families of extrapolation pairs following the convention in [Book_Extrapolation]. A family of extrapolation pairs $\mathcal{F}$ is a collection of pairs $(F,G)$ of measurable functions taking values in $\mathcal{K}_{\mathrm{bcs}}$ such that neither $F$ nor $G$ is equal to $\{0\}$ almost everywhere. In addition, given such a family $\mathcal{F},$ we call each element $(F,G)\in\mathcal{F}$ an extrapolation pair. We are interested in inequalities of the form
@@
in the sense that this holds for all pairs $(F,G)\in\mathcal{F}$ for which the left-hand side is finite and with the constant $C$ depending on the characteristic $[W]_{A_{p}}$ but not on the particular weight $W.$
Theorem A (Bownik, Cruz-Uribe [Cruz_Uribe_Bownik_Extrapolation]).
@@
The proof of this result is exactly the same as in [Cruz_Uribe_Bownik_Extrapolation, Theorem 9.1], up to the point that the calculations are done directly with the given convex body valued functions $F$ and $G,$ instead of taking $F(x)=\{\lambda f(x)\colon|\lambda|\leq 1\}$ and $G(x)=\{\lambda g(x)\colon|\lambda|\leq 1\}$ for given vector valued functions $f$ and $g$ (also note that in [Cruz_Uribe_Bownik_Extrapolation] $F(x)$ and $G(x)$ are segments with endpoints ±$\pm f(x)$ and ±$\pm g(x)$ because in that article real symmetric convex bodies are used, instead of complex symmetric). Since the proof in [Cruz_Uribe_Bownik_Extrapolation] follows closely that of the scalar case, we also refer the reader to [Book_Extrapolation, Chapter I.3] for a detailed exposition on the topic in the classical setting.
@@
It is obvious that $K$ is symmetric. Moreover, a standard weak star compactness argument yields that $K$ is closed. Let us check that $K$ is convex. For any sequences $\{v_{m}\}_{m=1}^{\infty},$$\{u_{m}\}_{m=1}^{\infty}$ with $v_{m},u_{m}\in K_{m},$ for all $m=1,2,\ldots,$ any complex numbers $a_{1},a_{2},\ldots$ and $b_{1},b_{2},\ldots$ with $\left\|{\{a_{m}\}}\right\|_{{q^{\prime}}}\leq 1$ and $\left\|{\{b_{m}\}}\right\|_{{q^{\prime}}}\leq 1,$ and any $t\in(0,1),$ we can write
$t\sum_{m=1}^{\infty}a_{m}v_{m}+(1-t)\sum_{m=1}^{\infty}b_{m}u_{m}$
$=\sum_{m=1}^{\infty}(t|a_{m}|+(1-t)|b_{m}|)$
$\cdot\left(\frac{t|a_{m}|}{t|a_{m}|+(1-t)|b_{m}|}\cdot\frac{a_{m}}{|a_{m}|}v_{m}+\frac{(1-t)|b_{m}|}{t|a_{m}|+(1-t)|b_{m}|}\cdot\frac{b_{m}}{|b_{m}|}u_{m}\right),$
with the obvious modifications if $a_{m}=0$ or $b_{m}=0$ for some $m\geq 1,$ thus it is clear that $t\sum_{m=1}^{\infty}a_{m}v_{m}+(1-t)\sum_{m=1}^{\infty}b_{m}u_{m}\in K.$
@@
Fix first $1<q<\infty.$ We construct a new family $\mathcal{F}_{q}$ of convex body valued functions as follows. For each sequence $\{(F_{n},G_{n})\}\subseteq\mathcal{F}$ such that
$\sum_{n=1}^{\infty}|F_{n}(x)|^{q}<\infty\quad\text{ and }\quad\sum_{n=1}^{\infty}|G_{n}(x)|^{q}<\infty$ (3.4)
@@
$F(x)=\Sigma_{q}(\{F_{n}(x)\}),\qquad G(x)=\Sigma_{q}(\{G_{n}(x)\}).$
Observe that by lemma 6, both $F$ and $G$ are convex body valued functions with
$|W(x)^{1/p}F(x)|\sim_{d,p}\left(\sum_{n=1}^{\infty}|W(x)^{1/p}F_{n}(x)|^{q}\right)^{1/q}$

### v3

@@ 2.3 Matrix $A_{p}$ weights
A function $W$ on $\mathbb{R}^{n}$ is said to be a $d\times d$-matrix valued weight, or just a matrix weight, if it is a locally integrable $\mathrm{M}_{d}(\mathbb{C})$-valued function such that $W(x)$ is a positive-definite matrix for a.e. $x\in\mathbb{R}^{n}.$ Given $1<p<\infty$ we define the norm
$\|F\|_{\mathrm{L}^{p}(W)}\coloneq\||W^{1/p}F|\|_{\mathrm{L}^{p}}$
for all $\mathrm{M}_{d}(\mathbb{C})$-valued measurable functions $F$ on $\mathbb{R}^{n}.$ For $p=\infty,$ this definition needs a slight modification. Namely, we define the norm
$\|F\|_{\mathrm{L}^{\infty}(W)}\coloneq\||WF|\|_{\mathrm{L}^{\infty}}$
for the same class of matrix valued functions as before.
@@
In this work we will be considering in particular one-parameter, respectively biparameter ($d\times d$) matrix valued $A_{p}$ weights on $\mathbb{R}^{n},$ respectively on $\mathbb{R}^{n}\times\mathbb{R}^{m},$ for $1<p<\infty.$ Following [36], we say that a matrix weight on $\mathbb{R}^{n}$ is a one-parameter $A_{p}$ weight if
$\sup_{R}\frac{1}{|R|}\int_{R}\left(\frac{1}{|R|}\int_{R}|W(x)^{1/p}W(y)^{-1/p}|^{p^{\prime}}\,\mathrm{d}y\right)^{p/p^{\prime}}\,\mathrm{d}x<\infty,$ (2.3)
@@
where the supremum ranges over all bounded cubes $R$ with sides parallel to the axes. We say that a matrix weight on $\mathbb{R}^{n}\times\mathbb{R}^{m}$ is a biparameter $A_{p}$ weight if the same condition holds when the supremum ranges over rectangles $R$ which are cartesian products of cubes on $\mathbb{R}^{n}$ and on $\mathbb{R}^{m}$ with sides parallel to the axes. For $p=1,$ note that condition (2.3) becomes
$\sup_{R}\operatorname*{ess\,sup}_{x\in R}\frac{1}{|R|}\int_{R}|W(x)^{-1}W(y)|\,\mathrm{d}y<\infty,$
with the same considerations as before distinguishing the one-parameter and biparameter cases. For $p=\infty,$ following [5], we say that a matrix weight is an $A_{\infty}$ weight if
$\sup_{R}\operatorname*{ess\,sup}_{x\in R}\frac{1}{|R|}\int_{R}|W(x)W(y)^{-1}|\,\mathrm{d}y<\infty,$
with identical considerations distinguishing the one-parameter and biparameter cases. For a more detailed explanation for these definitions and their relation to $A_{p}$ norm functions, see [5, Section 6]. Also, for a general overview on the topics of matrix weights, convex body valued functions, matrix extrapolation and their relation to the scalar theory, we refer the reader to [11].
In the second part of this work, dyadic versions of these weights adapted to dyadic grids in $\mathbb{R}^{n},$ respectively product dyadic grids in $\mathbb{R}^{n}\times\mathbb{R}^{m},$ will also be of major importance. Since the relevant definitions are by now standard, we omit repeating them and instead refer the reader to the detailed exposition in [14, Subsection 3.3]. Let us only stress for later convenience the following estimates from [14]. Let $\bm{\mathcal{D}}=\mathcal{D}^{1}\times\mathcal{D}^{2}$ be a product dyadic grid in $\mathbb{R}^{n}\times\mathbb{R}^{m}$ and let $W$ be a $(d\times d)$ matrix $\bm{\mathcal{D}}$-dyadic biparameter $A_{p}$ weight on $\mathbb{R}^{n}\times\mathbb{R}^{m}.$ Then
$W_{x_{1}}]_{A_{p},\mathcal{D}^{2}}\lesssim_{p,d}[W]_{A_{p},\bm{\mathcal{D}}},\quad\text{for a.e. }x_{1}\in\mathbb{R}^{n}.$ (2.4)
Moreover, if $Q$ is any cube in $\mathcal{D}^{2},$ then
$W_{Q}]_{A_{p},\mathcal{D}^{1}}\lesssim_{p,d}[W]_{A_{p},\bm{\mathcal{D}}}.$ (2.5)
@@ 2.4 Convex body valued functions
Here we consider functions taking values in the collection of closed bounded symmetric convex sets of $\mathbb{C}^{d}.$ Take into account that by symmetric here we mean complex symmetric. That is, a set $A\subseteq\mathbb{C}^{d}$ is complex symmetric (or just symmetric in this article) if for every $u\in A$ and every $\lambda\in\mathbb{C}$ with $|\lambda|=1$ it is also the case that $\lambda u\in A.$ Of course, if the set $A$ is convex in addition to symmetric, it will also be the case that, for every $u\in A$ and every $\lambda\in\mathbb{C}$ with $|\lambda|\leq 1,$ also $\lambda u\in A.$ In other words, the “symmetric convex sets” as we have defined them are precisely the balanced convex sets. We will denote the set of closed subsets of $\mathbb{C}^{d}$ by $\mathcal{K}(\mathbb{C}^{d}),$ or just $\mathcal{K}$ when the dimension of the ambient space is clear by the context. In addition, we define a convex body to be a closed bounded convex and symmetric subset of $\mathbb{C}^{d}.$ We will use the symbol $\mathcal{K}_{\mathrm{bcs}}(\mathbb{C}^{d})$ to denote the set of convex bodies on $\mathbb{C}^{d}$ and, whenever the dimension of the ambient space is unambiguous, simply by $\mathcal{K}_{\mathrm{bcs}}.$ Recall that if $K\in\mathcal{K}_{\mathrm{bcs}}(\mathbb{C}^{d})$ has nonempty interior, then there exists a unique ellipsoid $E\subseteq K$ with maximal volume. This ellipsoid is called the John ellipsoid and it satisfies
$E\subseteq K\subseteq\sqrt{d}E.$
We focus now on functions $F\colon\mathbb{R}^{n}\rightarrow\mathcal{K}_{\mathrm{bcs}}(\mathbb{C}^{d})$ and we gather the definitions and basic properties that we will need for our results. Some of these definitions can be found in more general forms in the texts that we cite; we will restrict though to the cases that are of interest to us to avoid an excess of concepts. For an introduction to such functions, their properties and how to define their integrals see [5, 12] and for a detailed exposition see [1].
Since throughout the article we will consider both functions taking values in $\mathbb{C}^{d}$ and functions taking values in $\mathcal{K}_{\mathrm{bcs}},$ we will use a typographic convention to avoid confusion. We will denote functions taking values in $\mathbb{C}^{d}$ with lowercase letters $f,g,h,\ldots$ On the other hand, we will use uppercase letters $F,G,H,\ldots$ to denote functions taking values in $\mathcal{K}_{\mathrm{bcs}}(\mathbb{C}^{d}).$ In any case, we will also explicitly state the target space of the functions we use.
@@
for every $x\in\mathbb{R}^{n}$ (see [1, Theorem 8.1.4]). The functions that form such sequences will be called selection functions for $F.$ In addition, the set of all selection functions for $F$ is denoted by $S^{0}(F).$ Observe that this is the same as the set of all measurable functions $f$ such that $f(x)\in F(x)$ for every $x\in\mathbb{R}^{n}.$ Here we restrict to functions taking values on $\mathcal{K}_{\mathrm{bcs}}(\mathbb{C}^{d})$ due to the applications that we consider later. Nonetheless, the previous definitions and concepts apply verbatim to functions taking values on $\mathcal{K}(\mathbb{C}^{d}).$ When one is interested in the norm of $F,$ it is possible to restrict to selection functions, as the following lemma shows (see [5, Lemma 3.9]). We include its proof for completeness, although the arguments of Bownik and Cruz-Uribe are equally valid for complex convex bodies in this case.
@@
$F_{0}(x)=\{v\in F(x)\colon|v|=g_{0}(x)\}=F(x)\cap(g_{0}(x)\mathbf{{S}}_{\mathbb{C}^{{d}}})$
taking values on $\mathcal{K}(\mathbb{C}^{d}),$ where $\mathbf{{S}}_{\mathbb{C}^{{d}}}=\{v\in\mathbb{C}^{d}\colon|v|=1\}$ denotes the complex (d1)(d-1)-dimensional sphere. Since both $F$ and $g_{0}\mathbf{{S}}_{\mathbb{C}^{{d}}}$ are measurable, so is $F_{0}$ (see [1, Theorems 8.2.2 and 8.2.4]).
@@
Given a $(d\times d)$-matrix weight $W,$ we define the weighted Lebesgue space $\mathrm{L}^{p}(W)$ of convex body valued functions as the set of functions $F$ for which $W(x)^{1/p}F(x)\in\mathrm{L}^{p}$ if $1\leq p<\infty$ and $W(x)F(x)\in\mathrm{L}^{\infty}$ if $p=\infty$. Finally, we denote $\|F\|_{\mathrm{L}^{p}(W)}:=\|W^{1/p}F\|_{\mathrm{L}^{p}}$ if $1\leq p<\infty$ and $\|F\|_{\mathrm{L}^{\infty}(W)}:=\|WF\|_{\mathrm{L}^{\infty}}$ if $1\leq p<\infty$.
@@ 3 Vector valued extensions of operators on convex body valued functions
The aim of this section is to prove vector valued estimates for convex body valued functions with matrix weights, as well as to apply it to vector valued estimates for the matrix weighted maximal operators for vector valued functions (see Subsection 3.1 for the definitions and details). This will be a consequence of an extrapolation result on spaces of functions with values on $\mathcal{K}_{\mathrm{bcs}}$ and it is a modification of the analogous result for matrix weights and vector valued functions due to Bownik and Cruz-Uribe (see [5, Section 9]). Before stating the extrapolation theorem, we introduce the concept of families of extrapolation pairs following the convention in [10]. A family of extrapolation pairs $\mathcal{F}$ is a collection of pairs $(F,G)$ of measurable functions taking values in $\mathcal{K}_{\mathrm{bcs}}$ such that neither $F$ nor $G$ is equal to $\{0\}$ almost everywhere. In addition, given such a family $\mathcal{F},$ we call each element $(F,G)\in\mathcal{F}$ an extrapolation pair. We are interested in inequalities of the form
@@
in the sense that this holds for all pairs $(F,G)\in\mathcal{F}$ for which the left-hand side is finite and with the constant $C$ depending on the characteristic $W]_{A_{p}}$ but not on the particular weight $W.$
Theorem A (Bownik, Cruz-Uribe [5]).
@@
This statement is the complex variable version of [5, Theorem 9.1]. Also, observe that [5, Theorem 9.1] refers to real vector valued functions, but Theorem A is stated directly for complex convex body valued functions. The reason is that, in [5], given real vector valued functions $f$ and $g$ Bownik and Cruz-Uribe define auxiliary real convex body valued functions $F$ and $G$ by taking the segments with endpoints $\pm f$ and $\pm g$ respectively, and the whole proof is done for these convex body valued functions. For the sake of generality, we opted to use convex body valued functions since the beginning. The proof of Theorem A follows that of [5, Theorem 9.1] almost verbatim and we do not include it here. However, there are some minor differences and a couple of technical details that require some care. The main difficulty in translating the real variable proof to the complex variable one is the following. In [5] Bownik and Cruz-Uribe consider a family of norms on $\mathbb{R}^{d}$ parametrized in a measurable way, such as the one given by a real (almost everywhere positive definite) matrix weight, and they include a proof of the fact that one can choose a family of reducing operators for such a family of norms. The proof of the analogous fact in the complex setting appears already in [14, Section 9]. We defer the discussion on Theorem A to Appendix A. There we review the proof of [5, Theorem 9.1]. Most of the steps are just commented briefly together with an explanation of why they hold all the same in the complex setting. We devote more detail though to those steps that are not immediate, namely the measurability of reducing operators. We also refer the reader to [10, Chapter I.3] for a detailed exposition on the topic in the classical setting.
@@
It is obvious that $K$ is symmetric. In addition, a direct computation shows that $K$ is convex.
Let us now show that $K$ is closed. In fact, we will show directly that $K$ is compact. This follows from a standard weak star compactness argument, whose details we present for the reader’s convenience.
Since $\mathbb{C}^{d}$ is a separable metric space, it suffices to show that $K$ is sequentially compact. Let $\{x_{r}\}^{\infty}_{r=1}$ be a sequence of elements of $K$. Then, for all $r=1,2,\ldots$ we can write
$x_{r}=\sum_{m=1}^{\infty}a^{r}_{m}v_{m}^{r},$
where $a^{r}=\{a^{r}_{m}\}_{m=1}^{\infty}\in\ell^{q^{\prime}}$, $\|a^{r}\|_{\ell^{q^{\prime}}}\leq 1$, and $v_{m}^{r}\in K_{m}$, for all $m=1,2,\ldots$.
It is well-known that $\ell^{q^{\prime}}$ is isometrically isomorphic to the dual $(\ell^{q})^{*}$ of $\ell^{q}$ under the natural pairing, since $q<\infty$ (see for instance [9, Chapter III, Examples 5.9 and 5.10]). Moreover, it is clear that $\ell^{q}$ is a a separable Banach space. Therefore, by the Banach–Alaoglou theorem ([9, Chapter V, Theorems 3.1 and 5.1]) we deduce that the closed unit ball of $\ell^{q^{\prime}}$ is sequentially compact under the weak star topology on $\ell^{q^{\prime}}$. Clearly, $\{a^{r}\}_{r=1}^{\infty}$ is a sequence in the closed unit ball of $\ell^{q^{\prime}}$. We deduce that after passing to a subsequence of $\{a^{r}\}^{\infty}_{r=1}$, one can find $a=\{a_{m}\}_{m=1}^{\infty}\in\ell^{q^{\prime}}$, $\|a\|_{\ell^{q^{\prime}}}\leq 1$ with $a^{r}\to a$ in the weak star topology of $\ell^{q^{\prime}}$ as $r\to\infty$. Notice that to ease notation, we have not introduced a new notation for this subsequence of $\{a^{r}\}_{r=1}^{\infty}$.
Notice that
$\|\{|a^{r}_{m}-a_{m}|\}_{m=1}^{\infty}\|_{\ell^{q^{\prime}}}\leq 2,\quad\forall r=1,2,\ldots.$
Clearly, a dilation of a compact set in a topological vector space remains a compact set. Thus, by the same token as before, after passing to a further subsequence we can find $b\in\ell^{q^{\prime}}$ with $\|b\|_{\ell^{q^{\prime}}}\leq 2$ such that
$\{|a^{r}_{m}-a_{m}|\}_{m=1}^{\infty}\to b$
in the weak star topology of $\ell^{q^{\prime}}$ as $r\to\infty$. Since convergence in the weak star topology implies pointwise convergence on $\mathbb{N}$ we deduce
$a^{r}\to a$
as well as
$\{|a^{r}_{m}-a_{m}|\}_{m=1}^{\infty}\to b$
pointwise on $\mathbb{N}$ as $r\to\infty$, showing that $b$ must be the zero sequence.
We now claim that
$\sum_{m=1}^{\infty}a^{r}_{m}v_{m}^{r}-\sum_{m=1}^{\infty}a_{m}v_{m}^{r}\to 0$ (3.4)
in $\mathbb{C}^{d}$ as $r\to\infty$. Indeed, we compute
$\left|\sum_{m=1}^{\infty}a^{r}_{m}v_{m}^{r}-\sum_{m=1}^{\infty}a_{m}v_{m}^{r}\right|\leq\sum_{m=1}^{\infty}|a^{r}_{m}-a_{m}|\cdot|v_{m}^{r}|\leq\sum_{m=1}^{\infty}|a^{r}_{m}-a_{m}|\cdot|K_{m}|.$
Observe that $\{|K_{m}|\}_{m=1}^{\infty}$ is a fixed sequence in $\ell^{q}$. Therefore, the weak star convergence of $\{|a^{r}_{m}-a_{m}|\}_{m=1}^{\infty}$ against the zero sequence in $\ell^{q^{\prime}}$ proves (3.4).
Now, for each $m=1,2,\ldots$, $K_{m}$ is a compact metric space, therefore the cartesian product $\prod_{m=1}^{\infty}K_{m}$ can also be made into a compact metric space. Thus, after passing to a further subsequence, one can find a sequence $\{v_{m}\}_{m=1}^{\infty}\in\prod_{m=1}^{\infty}K_{m}$ such that
$v_{m}^{r}\to v_{m}$
in $K_{m}$, that is in $\mathbb{C}^{d}$ as $r\to\infty$.
Lastly, we show that
$\sum_{m=1}^{\infty}a_{m}v_{m}^{r}\to\sum_{m=1}^{\infty}a_{m}v_{m}$ (3.5)
in $\mathbb{C}^{d}$ as $r\to\infty$. This will show that the sequence $\{x_{r}\}^{\infty}_{r=1}$ with which we began has a subsequence that converges to the element $x:=\sum_{m=1}^{\infty}a_{m}v_{m}$ of $K$.
Consider the counting measure $\mu$ on $\mathbb{N}$. Observe that
$\{a_{m}v_{m}^{r}\}_{m=1}^{\infty}\to\{a_{m}v_{m}\}_{m=1}^{\infty}$
pointwise on $\mathbb{N}$. Furthermore, notice that
$|a_{m}v_{m}^{r}|\leq|a_{m}|\cdot|K_{m}|,\quad\forall r=1,2,\ldots$
and that $\{|a_{m}|\cdot|K_{m}|\}_{m=1}^{\infty}\in L^{1}(\mathbb{N},\mu)$. Therefore, an application of the Dominated Convergence Theorem yields immediately (3.5).
@@
Fix first $1<q<\infty.$ We construct a new family $\mathcal{F}_{q}$ of convex body valued functions as follows. For each sequence $\{(F_{m},G_{m})\}^{\infty}_{m=1}\subseteq\mathcal{F}$ such that
$\sum_{m=1}^{\infty}|F_{m}(x)|^{q}<\infty\quad\text{ and }\quad\sum_{m=1}^{\infty}|G_{m}(x)|^{q}<\infty$ (3.6)
@@
$F(x)=\Sigma_{q}(\{F_{m}(x)\}^{\infty}_{m=1}),\qquad G(x)=\Sigma_{q}(\{G_{m}(x)\}^{\infty}_{m=1}).$
Observe that by Lemma 5, both $F$ and $G$ are convex body valued functions, $F,G:\mathbb{R}^{n}\to\mathcal{K}_{\mathrm{bcs}}$ with
$|W(x)^{1/p}F(x)|\sim_{d,p}\left(\sum_{m=1}^{\infty}|W(x)^{1/p}F_{m}(x)|^{q}\right)^{1/q}$